OpenStudy (comrad):

a 3.0 kg mass moving to the right at 1.4 m/s collides in a perfectly inelastic collision with a 2.0 kg mass initially at rest. what will the vleocity of the combined mass be after collision? show your work.

OpenStudy (comrad):

@airjordan200

OpenStudy (whpalmer4):

Do you have any idea how to work this problem?

OpenStudy (comrad):

not really

OpenStudy (whpalmer4):

The total momentum of the moving mass (before collision) and the stationary mass will be equal to the momentum of the combined mass (after collision).

OpenStudy (comrad):

?

OpenStudy (whpalmer4):

What is the momentum of the moving mass before the collision?

OpenStudy (comrad):

1.4 m/s

OpenStudy (whpalmer4):

No, that's the velocity. What is the momentum?

OpenStudy (comrad):

3.0 kg to the right?

OpenStudy (comrad):

im really horrible at physics. math and history are my strong points

OpenStudy (whpalmer4):

if you're good at math, physics should be fun :-) \[p=mv\]\(p\) is momentum, \(m\) is mass, \(v\) is velocity

OpenStudy (whpalmer4):

\(3.0 \text{ kg}\) is the mass moving to the right, with velocity \(1.4 \text { m/s}\) momentum of the moving object (before collision) is thus \[p=mv = (3.0\text{ kg})(1.4\text{ m/s}) = \]

OpenStudy (comrad):

so 4.2 m/s?

OpenStudy (whpalmer4):

No, did you keep all of the stuff you multiplied together? What happened to the \(\text{kg}\)?

OpenStudy (comrad):

4.2kg?

OpenStudy (whpalmer4):

treat units exactly like algebraic variables

OpenStudy (comrad):

4.2kg m/s lol

OpenStudy (whpalmer4):

bingo! okay, what about the momentum of the stationary object? What is that?

OpenStudy (comrad):

2.0kg with no velocity

OpenStudy (whpalmer4):

well, okay, the mass is \(2.0 \text{ kg}\) and velocity is \(0 \text{ m/s}\), what is the momentum? \[p=mv\]

OpenStudy (comrad):

2.0kg m/s so 4.2kg m/s *2.0kg m/s?

OpenStudy (comrad):

8.4kg m/s

OpenStudy (whpalmer4):

\[m =2.0\text{ kg}\]\[v = 0\text{ m/s}\]\[\text{momentum} = p = mv = \]

OpenStudy (comrad):

sso i was wrong?

OpenStudy (whpalmer4):

yes, in many ways :-) For starters, \[2.0 \text{ kg m/s} * 4.2\text{ kg m/s} = 2.0*4.2*\text{kg}*\text{kg}*\text{m}*\text{m}/(\text{s}*\text{s}) = 8.4 \text{kg}^2\text{m}^2/\text{s}^2\]

OpenStudy (whpalmer4):

remember, treat units just like you would \(x\) when doing algebra.

OpenStudy (comrad):

oooh o forgot the units that explains a lot of stuff

OpenStudy (whpalmer4):

\[4x*2x=8x^2\]

OpenStudy (comrad):

so is 8.4kg^2 m^2/s^2 the answer?

OpenStudy (whpalmer4):

and if you take the units along through the calculation, they provide a valuable error check on unit conversions, formula setup, etc. For example, the units of momentum have the dimension [mass][length]/[time], for example, \(\text{kg m}/\text{s}\) but your answer has different dimensions, does it not? That is a big red flag that something is rotten in Denmark, so to speak.

OpenStudy (whpalmer4):

As I said, momentum (\(p\)) is given by \[p = mv\]and we have \(m = 2.0 \text{ kg}\) and \(v = 0\text{ m/s}\) so the momentum of the stationary object (before the collision) is \[p = m v = (2.0\text{ kg})(0 \text{ m/s}) = \]

OpenStudy (comrad):

so what is the bloody answer then?

OpenStudy (comrad):

im more confused than when i first asked

OpenStudy (inkyvoyd):

lol don't worry about the answer, worry about knowing how to get there

OpenStudy (inkyvoyd):

most of the time answer is only worth 2 of 10 points anyways...

OpenStudy (whpalmer4):

I'm trying to get you to do it :-) Look, you find the momentum of all of the pieces before the collision. The momentum of an object is simply the mass * velocity. We found the momentum of the moving object. Now you need to find the momentum of the stationary object. I did everything except punch the buttons for you... \[p = mv = (2.0\text{ kg})(0 \text{ m/s}) = \]

OpenStudy (whpalmer4):

You know how to multiply by 0, don't you?

OpenStudy (comrad):

the question is worth 3 points

OpenStudy (comrad):

so the answer is 0? oh my god, this is confusing!

OpenStudy (inkyvoyd):

@Comrad , @whpalmer4 is trying to guide you thru the process of learning how to do this problem... You seem to be a little too eager to get the answer.

OpenStudy (whpalmer4):

Yes, the momentum of the stationary object is 0. Objects at rest have no momentum. So, the total momentum of the system before collision is the sum of the momentums of the moving object and the stationary object. \[4.2\text{ kg m/s} + 0 \text{ kg m/s} = \]

OpenStudy (comrad):

0kg m2/s^2

OpenStudy (whpalmer4):

no.... :-( Let's try something else. \[4.2x + 0x = \]

OpenStudy (comrad):

0x so 0 there is no value that could make the answer something else other than 0.

OpenStudy (comrad):

no value for x should i say

OpenStudy (whpalmer4):

No, \[4.2x + 0x = (4.2+0) x = 4.2x\]!!!

OpenStudy (whpalmer4):

4.2 miles + 0 miles = how many miles? 4.2 miles, or 0 miles?

OpenStudy (comrad):

my bad. the plus came through as a multiplication symbol. sorry!!!!

OpenStudy (comrad):

4.2 miles

OpenStudy (comrad):

*facepalm*

OpenStudy (whpalmer4):

Okay, so our total momentum is \[4.2 \text{ kg m/s} + 0 \text{ kg m/s} = 4.2 \text{ kg m/s}\] And after the collision, the two objects are now one object, with that same momentum, and the mass of the object is \[3.0\text{ kg}+2.0\text{ kg} = 5.0\text{ kg}\]

OpenStudy (whpalmer4):

Using \[p = mv\]find \(v\) such that \(p = mv\) is true when \(m = 5.0\text{ kg}\) and \(p = 4.2 \text{ kg m/s}\)

OpenStudy (whpalmer4):

\[p = mv\]\[4.2 \text{ kg m/s} = (5.0\text{ kg}) v\]

OpenStudy (comrad):

.84kg m/s^2

OpenStudy (whpalmer4):

You got the number right, the units are wrong, which means the answer is wrong. \[4.2 \text{ kg m/s} = (5.0 \text{ kg}) v\] \[v = \frac{4.2}{5.0} \frac{\text{ kg m/s}}{\text{kg}} = 0.84 \frac{\cancel{\text{ kg}}\text{ m/s}}{\cancel{\text{kg}}} = 0.84 \text{ m/s}\]

OpenStudy (comrad):

k

OpenStudy (comrad):

so it only square when multiplying 2 velocities?

OpenStudy (comrad):

*squared

OpenStudy (whpalmer4):

Here's what we did. We had two objects, with masses \(m_1 = 3.0 \text{ kg}\) and \(m_2 = 2.0 \text{ kg}\) and velocities before collision of \(v_1 = 1.4\text{ m/s}\) and \(v_2 = 0\text{ m/s}\) momentum is conserved in an inelastic collision, so: \[p_1 + p_2 = p_{final}\]\[m_1*v_1 + m_2*v_2 = (m_1+m_2)v_{final}\] We just plugged in the numbers and solved for \(v_{final}\) Yes. Treat units exactly like you would algebraic variables. \[2x+3x = 5x\]\[2x*3x=6x^2\]

OpenStudy (whpalmer4):

Because our \(v_2=0\), our equation simplified to \[m_1*v_1 = (m_1+m_2)v_{final}\] You can see that if the mass of the stationary object is much larger, the collision will essentially result in the first object stopping and not much else. Think about firing a rifle bullet into the ground. The Earth doesn't exactly get knocked out of its orbit :-) If the situation is reversed, where a bit object hits a small one, think dump truck hitting some poor schmuck on his bicycle, the big object continues along pretty much at the same speed, with a new hood ornament. Ouch.

OpenStudy (comrad):

ok so can you put all of our equations and answers on 1 things? its kinda confusing

OpenStudy (comrad):

i hate physics

OpenStudy (whpalmer4):

Okay. Moving object: \[m_1 = 3.0 \text{ kg}\]\[v_1 = 1.4\text{ m/s}\] Stationary object: \[m_2 = 2.0 \text{ kg}\]\[v_2 = 0\text{ m/s}\] Combined objects after collision: \[m_{final} = m_1+m_2\] \[p_1 + p_2 = p_{final}\]\[m_1 * v_1 + m_2*v_2 = (m_1+m_2)v_{final}\]\[v_{final} = \frac{m_1*v_1 + m_2*v_2}{m_1+m_2}\]\[v_{final} = \frac{(3.0\text{ kg})(1.4\text{ m/s}) + (2.0\text{ kg})(0\text{ m/s})}{(3.0\text{ kg}+2.0\text{ kg})} = \frac{4.2\text{ kg m/s}}{5.0\text{ kg}}\]\[v_{final} = 0.84\text{ m/s}\]

OpenStudy (comrad):

so thats the final answer? this is just the most confusing thing ever? i mean it makes since mathematicaly but i still dont get it

OpenStudy (comrad):

There ya go!

OpenStudy (whpalmer4):

"what will the vleocity of the combined mass be after collision?" the velocity of the combined mass after collision will be 0.84 m/s.

OpenStudy (comrad):

thanks man. ur a life saver.

OpenStudy (whpalmer4):

We roughly double the mass, so the velocity is roughly cut in half.

OpenStudy (comrad):

wait, so this isn't the answer?

OpenStudy (whpalmer4):

No, you have the answer, I'm just trying to help you get a feel for what's going on.

OpenStudy (comrad):

ooooh ok.

OpenStudy (whpalmer4):

Momentum before = momentum after. That's all there is to it.

OpenStudy (comrad):

Well thanks!

OpenStudy (whpalmer4):

you're welcome! time for me to get back to my real work...